OpenStudy (linn99123):

What are the solutions to the equation x2 + 6 = 87? A. –6 and 6 B. –10 and 10 C. –9 and 9 D. The equation has no solutions. im thinking its A

OpenStudy (anonymous):

Simplifying x2 + 6 = 87 Reorder the terms: 6 + x2 = 87 Solving 6 + x2 = 87 Move all terms containing x to the left, all other terms to the right. Add '-6' to each side of the equation. 6 + -6 + x2 = 87 + -6 Combine like terms: 6 + -6 = 0 0 + x2 = 87 + -6 x2 = 87 + -6 Combine like terms: 87 + -6 = 81 x2 = 81 Simplifying x2 = 81 Take the square root of each side: x = {-9, 9}

OpenStudy (anonymous):

So therefore your answer would be C!

OpenStudy (linn99123):

o dang i was way off :O

OpenStudy (anonymous):

Haha, it's okay we all make mistakes!

OpenStudy (anonymous):

Glad I could help! See ya!